# WolBanking77: Wolof Banking Speech Intent Classification Dataset

### v2

@@ 2.2 Summary of contributions
  • The audio dataset contains 263 sentences covering 10 intents in the banking and transport domains, with diverse voices and accents, making it the first of its kind in the Wolof language at the time of writing.
@@ 5.3 Model training
All models presented in this article are based on Pytorch library [Paszke et al., 2019]. All experiments were run on Runpod  [^12] https://www.runpod.io/, and Kaggle. [^13] https://www.kaggle.com/ P100 GPU (16GB VRAM) was used to finetune BERT-Base, mDeBERTa-v3-base, Afro-xlmr-large and AfroLM_active_learning. RTX 4090 GPU for AfriteVa V2. For Zero-shot and few-shot classification, RTX 2000 Ada GPU (16 GB VRAM) was used. Few-shot classification is done with SetFit huggingface library [Tunstall et al., 2022]. 2-shots refers to 2 samples per intent while 8-shots represents 8 samples per intent. All pre-trained text models are hosted on huggingface platform, they are fine-tuned using transformers library Wolf et al. [2020] except for Llama-3.2 which is fine-tuned using torchtune library [torchtune maintainers, 2024].
@@
Canary Flash and Phi-4-multimodal-instruct ASR models are fine-tuned on the Wolbanking77 audio dataset using the NVIDIA NEMO framework Kuchaiev et al. [2019] and Huggingface Transformers library, respectively, on an A100 SXM GPU (80GB VRAM). In contrast, Distil-whisper-large-v3.5 is fine-tuned using the Huggingface Transformers framework on an RTX 2000 Ada GPU (16 GB VRAM). All ASR models are fine-tuned for 1,000 steps.

### v3

@@ 2.2 Summary of contributions
  • An audio dataset with 263 sentences covering 10 intents in the banking and transport domains, including diverse voices and accents, making it the first of its kind in the Wolof language at the time of writing.
@@ 5.3 Model training
All models presented in this article are based on Pytorch library [Paszke et al., 2019]. All experiments were conducted on Runpod  [^12] https://www.runpod.io/ and Kaggle. [^13] https://www.kaggle.com/ P100 GPU (16GB VRAM) was used to finetune BERT-Base and mDeBERTa-v3-base. RTX 4090 GPU for Afro-xlmr-large, AfroLM_active_learning, AfriteVa V2 and Llama-3.2 models. For Zero-shot and Few-shot classification, RTX 2000 Ada GPU (16 GB VRAM) was used. Few-shot classification was performed using SetFit huggingface library [Tunstall et al., 2022]. 2-shots refers to 2 samples per intent while 8-shots represents 8 samples per intent. All pre-trained text models are hosted on huggingface platform, they were fine-tuned using transformers library Wolf et al. [2020] except for Llama-3.2 which was fine-tuned using torchtune library [torchtune maintainers, 2024].
@@
Canary Flash and Phi-4-multimodal-instruct ASR models were fine-tuned on the Wolbanking77 audio dataset using the NVIDIA NEMO framework Kuchaiev et al. [2019] and Huggingface Transformers library, respectively, on an A100 SXM GPU (80GB VRAM). In contrast, Distil-whisper-large-v3.5 was fine-tuned using the Huggingface Transformers framework on an RTX 2000 Ada GPU (16 GB VRAM). All ASR models were fine-tuned for 1,000 steps.